Metlock Inc. loans money to John Kruk Corporation in the amount of $944,000. Metlock accepts an 8% note due in 7 years with interest payable semiannually. After 2 years (and receipt of interest for 2 years), Metlock needs money and therefore sells the note to Chicago National Bank, which demands interest on the note of 10% compounded semiannually. What is the amount Metlock will receive on the sale of the note?
Answer #1
Amount received on the sale of the note 871104
Workings:
Seni-annual interest 37760 =944000*8%*6/12
X PV factor 7.72173 =(1-(1.05)^-10)/0.05
Present value of interest 291573
Principal 944000
X PV factor 0.61391 =1/1.05^10
Present value of principal 579531
Amount received on the sale of the note 871104 =291573+579531